OpenStudy (ilovebmth1234):

Widget Wonders produces widgets. they have found that the cost, C(X), of making x wigets is a quadratic function in terms of x. the company also discovers that it costs $30 to produce 2 widgets, $118 to produce 4 widgets, and $766 to produce 10 widgets. find the total cost of producing 6 widgets.

OpenStudy (ilovebmth1234):

@LastDayWork

OpenStudy (lastdaywork):

In general, the quadratic function can be written as \[C(x) = px^2+qx+r\] We are given that - C(2) = 30 C(4) = 118 C(10) = 766 Find the value of p, q, r Then calculate C(6)

OpenStudy (ilovebmth1234):

i forget how to do that?

OpenStudy (ilovebmth1234):

c(4)=a4^2+m4+b c(2)=a2^2+m2+b c(10)=a10^2+m10+b

OpenStudy (lastdaywork):

c(4)=a4^2+m4+b = 118 c(2)=a2^2+m2+b = 30 c(10)=a10^2+m10+b = 766 So we have three equations and three unknowns..Solve them...if you dare :P

OpenStudy (lastdaywork):

Sorry, I have to go now..I'll return in an hour Also, I cant PM you :(

OpenStudy (ilovebmth1234):

its alright thanks for the help today :)

OpenStudy (anonymous):

i'm sorry I don't know but I will keep trying!

OpenStudy (ilovebmth1234):

grrrr!!!

OpenStudy (anonymous):

let c(x) = ax^2 +bx + c 30 = 4a + 2b + c (1) 118 = 16a + 4b + c (2) 766 = 100a + 10b + c (3) (2) -(1) gives 88 = 12a + 2b (4) (3) -(2) gives 648 = 84a + 6b (5) (4) x 3 gives 264 = 36a + 6b (6) (5) - (6) gives 384 = 48a a = 384/48 = 192 /24 = 96/12 = 8 Sub. a = 8 into (4) 88 = 96 + 2b 2b = -8 b = -4 Sub. a=8 & b=-4 into (1) 30 = 4a + 2b + c (1) 30 = 32 -8 +c c = 30 -32 +8 = 6 so c(x) = 8x^2 - 4x + 6 I think this is it I hope so I tried my best!! if not I am sorry

OpenStudy (ilovebmth1234):

thanks

OpenStudy (anonymous):

you are welcome was that it?

OpenStudy (ilovebmth1234):

whats the total value though?

OpenStudy (anonymous):

okay hold on I will find that out too just give me a few sec's okay brb

OpenStudy (ilovebmth1234):

ok :)

OpenStudy (anonymous):

30 118 766 2 4 x 118 766 4 x 118x= 3064 x=3064/ 118 x=25.96 or 26 widgets the total is 25.96 or 26.00 I think through

OpenStudy (ilovebmth1234):

im supposed to find out how much it cost to make 6 widgets

OpenStudy (anonymous):

oh I am sorry let me try again

OpenStudy (anonymous):

Sub. a=8 & b=-4 into (1) 30 = 4a + 2b + c (1) 30 = 32 -8 +c c = 30 -32 +8 = 6 so c(x) = 8x^2 - 4x + 6 if you are looking for =6 this is the one that you will need. :)

OpenStudy (ilovebmth1234):

thank you and i know this is probably frustrating because it was for me

OpenStudy (anonymous):

yeah it is hard and I hate math so much and I cant wait to be done with math

OpenStudy (ilovebmth1234):

me either i have 3 more lessons and in done with mine for the year

OpenStudy (anonymous):

I think it would be $148

OpenStudy (ilovebmth1234):

it wasnt :/